Project

General

Profile

2015-11-17 memo-physics-report » History » Version 1

Rogers, Chris, 17 November 2015 12:52

1 1 Rogers, Chris
h1. 2015-11-17 memo-physics-report
2
3
2015/03
4
* Third draft run plan document is at http://micewww.pp.rl.ac.uk/projects/analysis/documents
5
* Lots of uncertainty around hardware, so we shall see what happens (ops meeting later)
6
* FC alignment:-
7
** Rogers has looked at settings, ~ 50 A solenoid mode looks okay (or 70 A flip mode); still not finalised
8
** Sophie is back from vacation and will start looking through "reconstruction" i.e. can we get good enough resolutions with TOF cut?
9
* Q789 switched off for TOF0 alignment
10
** nb this is a short run but should be done for TOF0 alignment
11
* SSU switched on with ECE:-
12
** Uncertainty around the hardware
13
** VB looking at trim coil currents
14
** Nb: needs some modelling of iron here - feeding in Jo Langlands (Sheffield grad student) to this effort + field mapping
15
* Straight tracks
16
** Staring at MuScat paper, looks like they were systematics dominated (errors ~ 50k muons)
17
** Working on the statistics analysis (can we get 50 k muons within momentum cut and radial cut at the absorber? what should the cuts be?)
18
19
3/4 December Review
20
* Step IV
21
** asked Ao to push through a new round of optics studies - he was distracted by PASI last week
22
** will then need to process at least one setting with full reconstruction
23
* Demonstration of ionisation cooling
24
** Baseline and baseline w/o M1 is at a reasonable point, using RMS emittance analysis
25
** CR is looking at generic phase space volume analyses (not RMS emittance); to fight the "do nothing" crowd
26
** Jaroslaw is looking at using third FC as a matching element (no news yet, Rogers pestering)
27
28
January optics Review
29
* v1 agenda handed to KL, v2 attached (some familiar names)
30
* await revised ToR - but we need to start the writing up process now to have a hope to converge by January; agenda becomes the latex skeleton
31
* I am scared by the timescale
32
33
Step IV
34
* Precision measurement of emittance
35
** VB working through the "TOF tracks" analysis (maybe she is in the meeting?), turning it into a reconstruction routine
36
** Awaits tracker reconstruction/geometry fixes
37
* Tracker alignment
38
** Discussion with Melissa/Jason/Durga/Tony Millington
39
** John Palin et al aligned trackers to monuments on the end face of SSU/SSD
40
** Tony Millington et al surveyed monuments on the outer rim of SSU/SSD
41
** end-face monuments are now obscured
42
** Jason Tarrant is in contact with John Palin to recover any (raw) survey data he may have, hand off to Tony
43
** Tony Millington will make a remedial survey of SSU on Wednesday this week when the diffuser goes in
44
** remedial survey on SSD may not be possible due to interference from cables and pipes
45
46
Cooling demo
47
* No update on the paper, chasing on Jaroslaw "cooling demo with focus coil" optics
48
* No update on the tolerances study, will chase Alan